5. A nonparametric procedure for two independent samples - A large sample case Aa Aa E Bullying, according to noted expertYou are concerned that the anxiety scores from bully-victims and bystanders may not satisfy the assumptions for a parametricThe Zobt for the rank sums statistic is -0.88 0.85 -1.81 - 50s that follow and decide whether to reject the null hypothesis,The or less than E decision is to the With a = .01, the region of rejection is defined as a z-score greater than Z-score for. The the With a = .01, the region of rejection is defined as a z-score greater than or less than Z-score for the rank sums s. The the With a = .01, the region of rejection is defined as a z-score greater than or less than Z-score for the rank sums s. The With a = .01, the region of rejection is defined as a z-score greater than or less than Z-score for the rank sums stati
Answer #1
sample size n1 group 1 48
sample size n2 group 2 52
sample mean = μ=n1(n1+n2+1)/2= 2424

You are concerned that the anxiety scores from bully-victims and bystanders may not satisfy the assumptions for a parametric

Variance= σ2=n1n2(n1+n2+1)/12= 21008.00
standard deviation= σ = 144.94
test stat z= Z=(U-μ)/σ = -1.81

. The With a = .01, the region of rejection is defined as a z-score greater than 2.58 Z-score for the rank sums statistic is